# Are EU low-carbon structural funds efficient in reducing emissions?

### v2

@@ 4.4 Effect by type of gases
We break down emissions by specific gases, considering CO2, CH4, and N2O, to assess the differentiated impact of low-carbon investments. We adopt the same methodological approach used for total GHG emissions: decomposing the per capita and intensity time series for each gas into trend and cycle components, and estimating the effects of expenditures on these indicators following the specification in Eq.(3). Results are reported in Tables 6 and 7, corresponding to the trend and cycle components, respectively. For readability, the estimated coefficients of control variables are omitted from the tables.
@@ 5 Discussion
When we employ instrumental variables to address potential endogeneity, the findings further reinforce the existence of regional disparities. The positive effect on per capita emissions in less developed regions is clearly confirmed, and additional significant effects emerge for transition regions that were not visible in the initial analysis. While this might suggest that estimations for transition regions were underestimated, the overall pattern is strengthened, reflecting heterogeneous regional impacts. Hence, the IV analysis enhances the credibility of a causal interpretation of the differentiated expenditure effects.
@@
Altogether, our findings provide strong empirical support for our central hypothesis. While we find consistent evidence that low-carbon economy funding does not lead to long-term reductions in emissions indicators in more developed regions, the adverse effects observed in less developed regions are particularly concerning, given that these areas are the primary targets of ESI fund support. Nevertheless, the results suggest that while the funds may not be highly effective in driving long-term decarbonisation, they could still play a stabilising role by dampening short-term emissions volatility.
@@
These findings call for a reassessment of EU climate investment strategies. On the one hand, the failure to curb long-term emissions in less developed areas raises questions about allocative efficiency and the risk of reinforcing carbon lock-in (Unruh, 2000). On the other hand, mechanisms behind the heterogeneous impacts – such as institutional capacity gaps, infrastructure rigidity, or lack of policy complementarities – remain underexplored and merit further attention (Rodríguez-Pose and Garcilazo, 2015; Charron et al., 2014). In this respect, Iammarino et al. (2019) emphasise the need for tailored policies that are sensitive to the structural prospects of different regions, aligning with the Barca report’s advocacy for place-based strategies (Barca, 2009).
@@ 6 Conclusions
This paper examines the climate impact of one of the European Union’s key instruments for regional development: the “Low-Carbon Economy” funds allocated through the European Structural and Investment funds during the 2007-2013 and 2014-2020 programming periods. Focusing on per capita emissions and emissions intensity – both decomposed into trend and cyclical components – we estimate the structural (long-run) and cyclical (short-run) effects of low-carbon expenditures across regions.
Our findings point to a substantial degree of heterogeneity. In more developed and transition regions, low-carbon funding shows no consistent long-term effect in reducing emissions, whether measured per capita or relative to GDP. However, such investments are associated with significant increases in long-term emissions trends in less developed regions, despite being the primary beneficiaries of these funds. These results suggest that, in lagging regions, climate-related investments may stimulate economic activity without achieving the intended decoupling from carbon-intensive processes. Instrumental variable estimations reinforce the robustness of our findings and lend support to a causal interpretation of the observed relationships.
@@ Acknowledgements
Low-carbon economy definitions by ESI programmes
@@
Estimation results modelled expenditures
The cohesion data platform presents estimations of the expenditure by NUTS 2, which are aggregated for the most significant types of structural funds only. Therefore, to address our need for yearly data, we have opted for two approaches: the one presented in the main text, which assumes that expenditures are distributed evenly over time (dividing the reported expenditures over the programming period), and the another one, presented here, which assumes they follow the same pattern as the total aggregated expenditure reported by the cohesion data platform.888Data available at: https://cohesiondata.ec.europa.eu/Other/Historic-EU-payments-regionalised-and-modelled/tc55-7ysv/about_data. Both strategies yielded qualitatively similar conclusions, as a robustness check the results of the second strategy are shown in Tables A2 and A3.
@@
Testing for instruments exogeneity
@@
| Trend residual | Cycle residual | |||
| (3t) | (6t) | (3c) | (6c) | |
| GHGEpc | GHGEi | GHGEpc | GHGEi | |
| Labour productivity ppet-8 | 0.003 | 0.002 | -0.002 | 0.000 |
@@
| R-squared | 0.000 | 0.001 | 0.001 | 0.000 |
| Notes: Standard errors are in parentheses. Significance level: ∗p0.1; ∗∗p0.05; ∗∗∗p0.01 | ||||

### v3

@@ 4.3 Effect by type of gases
We break down emissions by specific gases, considering CO2, CH4, and N2O, to assess the differentiated impact of low-carbon investments. We adopt the same methodological approach used for total GHG emissions: decomposing the per capita and intensity time series for each gas into trend and cycle components, and estimating the effects of expenditures on these indicators following the specification in Eq.(3). Results are reported in Tables 4 and 5, corresponding to the trend and cycle components, respectively. For readability, the estimated coefficients of control variables are omitted from the tables.
@@
Therefore, the gas-specific results further support H1, showing that the effects of low-carbon expenditures differ across levels of regional development and are also highly heterogeneous across emission types. At the same time, the evidence also appears to support H2, since in less-developed regions, the low-carbon expenditures are positively associated with CO2, CH4, and N2O trends, whereas in transition and developed regions, a negative relationship appears more common, although in many specifications this result is not statistically significant.
@@ 5 Discussion
Overall, the results are consistent with Hypothesis 1, indicating that low-carbon expenditures have heterogeneous impacts across different levels of regional development. Moreover, while we find consistent evidence that low-carbon economy funding does not lead to long-term reductions in emission indicators in more developed regions, the adverse effects observed in less developed areas are particularly concerning, given that these regions are the primary targets of ESI fund support. This pattern aligns with Hypothesis 2, which states that the effectiveness of such expenditures is conditional on regional development. Nevertheless, when examining the cyclical component of the emissions time series, we find that low-carbon investments may play a stabilising role by dampening short-term emissions volatility.
@@
These findings call for a reassessment of EU climate investment strategies. On the one hand, the failure to curb long-term emissions in less developed areas raises questions about allocative efficiency and the risk of reinforcing carbon lock-in (Unruh, 2000). On the other hand, mechanisms behind the heterogeneous impacts—such as institutional capacity gaps, infrastructure rigidity, or lack of policy complementarities—remain underexplored and merit further attention (Rodríguez-Pose and Garcilazo, 2015; Charron et al., 2014). In this respect, Iammarino et al. (2019) emphasise the need for tailored policies that are sensitive to the structural prospects of different regions, aligning with the Barca report’s advocacy for place-based strategies (Barca, 2009).
@@
It is worth noting that the identification strategy, based on long lags and fixed effects, is aimed to mitigate concerns of reverse causality; however, it may not fully address endogeneity. Instrumental variable estimations, using regional structural and institutional variables as instruments, resulted in weak instruments. Therefore, future work should seek to strengthen causal inference, for instance by exploiting exogenous quasi-natural experiments (e.g., changes in eligibility thresholds or policy allocation rules).
@@ 6 Conclusions
This paper examines the climate impact of one of the European Union’s key instruments for regional development: the “Low-Carbon Economy” funds allocated through the European Structural and Investment funds during the 2007-2013 and 2014-2020 programming periods. Focusing on per capita emissions and emissions intensity—both decomposed into trend and cyclical components—we estimate the structural (long-run) and cyclical (short-run) effects of low-carbon expenditures across regions.
Our findings point to a substantial degree of heterogeneity. In more developed and transition regions, low-carbon funding shows no consistent long-term effect in reducing emissions, whether measured per capita or relative to GDP. However, such investments are associated with significant increases in long-term emissions trends in less developed regions, despite being the primary beneficiaries of these funds. These results suggest that, in lagging regions, climate-related investments may stimulate economic activity without achieving the intended decoupling from carbon-intensive processes. These findings are consistent across specifications with different sets of controls.
@@
Appendix A1 Low-carbon economy definitions by ESI programmes
@@
Appendix A2 Estimation results modelled expenditures
The cohesion data platform presents estimations of the expenditure by NUTS 2, which are aggregated for the most significant types of structural funds only. Therefore, to address our need for yearly data, we have opted for two approaches: the one presented in the main text, which assumes that expenditures are distributed evenly over time (dividing the reported expenditures over the programming period), and the another one, presented here, which assumes they follow the same pattern as the total aggregated expenditure reported by the cohesion data platform.999Data available at: https://cohesiondata.ec.europa.eu/Other/Historic-EU-payments-regionalised-and-modelled/tc55-7ysv/about_data. Both strategies yielded qualitatively similar conclusions, as a robustness check the results of the second strategy are shown in Tables A2 and A3.
@@
Appendix A3 Instrumental variables (IV/2SLS)
An instrumental variables approach can be used to address potential endogeneity concerns related to low-carbon investments as a complement to our baseline fixed-effects specification. While the use of lagged explanatory variables in the baseline models helps mitigate simultaneity bias, endogeneity may persist due to omitted variables. For instance, more proactive regions might both attract higher low-carbon funding and implement complementary environmental measures not fully captured by our controls, potentially biasing the estimated coefficients.
However, one of the main challenges in identifying valid instruments for low-carbon investments is that, while several structural variables can plausibly explain the allocation of such funds without being directly correlated with emissions, the delayed nature of their effects substantially weakens the instruments’ explanatory power. As a result, the fitted values of investment exhibit limited variation, which reduces the likelihood of detecting statistically significant effects in the second stage. We present these results below.
We first identified a set of potential instruments for per capita low-carbon expenditures based on regional structural indicators from the ARDECO database, including labour productivity per employee (log), consumption of fixed capital (log), hours worked per capita, and compensation of employees (log). Then, we employed two-stage least squares estimations (2SLS), including year, regional, and eligibility fixed effects, clustering standard errors at the regional level, and controlling for lagged sectoral value added and structural indicators.
After testing various specifications, we found that, according to the underidentification test, these variables performed reasonably well when considered either individually or in combination. Subsequently, we observed that using three- to five-period lags yielded comparatively better first-stage results. It is not surprising that lagged versions of these variables perform well, as pre-existing structural conditions are likely to influence the allocation of funds. However, these lags are added to the five-period lag already applied to the dependent variable, weakening their potential as strong instruments.
In Table A4, we report the results for two specifications: the trend of GHG emissions per capita (third column) and the trend of GHG emissions intensity (fourth column), along with their corresponding first-stage regressions (first two columns). We present these estimations as an example to illustrate the limitations of our empirical approach. In these specifications, the set of instruments is reasonably correlated with the endogenous regressor, as indicated by the Kleibergen-Paap rk LM statistic (greater than 10, with Chi-sq. -value 0.05), and there is no evidence against their validity according to the Hansen J test (-value 0.1). However, the instruments consistently fail to pass the weak identification tests: the Kleibergen-Paap rk Wald statistic remains well below conventional thresholds (). Because this statistic is robust to heteroskedasticity and clustering, it provides the appropriate benchmark in our panel setting.
After implementing different specifications to exhaustively explore the possibility of addressing potential endogeneity concerns, we therefore conclude that the IV estimates cannot provide reliable evidence and must be interpreted with caution.
| First stage | Second stage | |||
@@
| Low Carbon Exp. pc | Low Carbon Exp. pc | GHGE per capita | GHGE intensity | |
| Low Carbon Exp. pc | -0.604 | -0.101 | ||
| (0.588) | (0.615) | |||
| Low Carbon Exp. pc | 3.071 | -7.180 | ||
| (2.064) | (4.764) | |||
| Low Carbon Exp. pcTER | 1.558* | 2.676** | ||
| (0.888) | (1.272) | |||
| Low Carbon Exp. pcTER | -3.332 | -3.195 | ||
| (4.136) | (5.033) | |||
| Low Carbon Exp. pcLDR | 0.535 | 0.446 | ||
| (0.693) | (0.726) | |||
| Low Carbon Exp. pcLDR | 2.078 | 7.318 | ||
| (2.252) | (4.569) | |||
| EPSt-5 | 0.018*** | 0.003** | -0.004 | -0.015 |
@@
| GVAO-U,t-6 | -0.128** | 0.015 | 0.112 | 0.233* |
| (0.052) | (0.014) | (0.109) | (0.137) | |
| Labour productivity ppet-8 | 0.238*** | 0.029* | ||
| (0.063) | (0.017) | |||
| Labour productivity ppet-8TER | -0.217*** | -0.024 | ||
| (0.066) | (0.021) | |||
| Labour productivity ppet-8LDR | -0.149** | -0.027 | ||
| (0.060) | (0.017) | |||
| Consumption of fixed capitalt-8 | -0.047 | -0.059*** | ||
| (0.034) | (0.013) | |||
| Consumption of fixed capitalt-8TER | 0.174*** | 0.046*** | ||
| (0.030) | (0.012) | |||
| Consumption of fixed capitalt-8LDR | 0.091*** | 0.009 | ||
| (0.025) | (0.010) | |||
| Hours worked per capitat-8 | 0.000 | 0.000 | ||
| (0.000) | (0.000) | |||
| Hours worked per capitat-8TER | -0.000 | 0.000 | ||
| (0.000) | (0.000) | |||
| Hours worked per capitat-8LDR | 0.000 | 0.000 | ||
| (0.000) | (0.000) | |||
| Compensation of Employeest-8 | 0.125*** | 0.011 | ||
| (0.037) | (0.016) | |||
| Compensation of Employeest-8TER | -0.047 | -0.013 | ||
| (0.058) | (0.024) | |||
| Compensation of Employeest-8LDR | -0.156*** | -0.005 | ||
| (0.052) | (0.021) | |||
| -1.384** | -0.315** | |||
| (0.633) | (0.157) | |||
@@
| Kleibergen-Paap rk LM (p-value) | 26.844 (0.000) | 26.844 (0.000) | ||
| Cragg–Donald Wald F | 12.957 | 12.957 | ||
| Kleibergen-Paap rk Wald F | 2.676 | 2.676 | ||
| Hansen J (p-value) | 7.542 (0.274) | 9.196 (0.163) | ||
| Notes: Robust standard errors clustered at the NUTS 2 level are reported in parentheses. Significance level: ∗p0.1; ∗∗p0.05; ∗∗∗p0.01. | ||||